28) find all solution between o through 2pi
2sin^2x= 2+cosx

Answer by stanbon(75887)

find all solution between o through 2pi
2sin^2x= 2+cosx
-----
Identity: sin^2(x) = 1-cos^2(x)
Substitute to get:
2(1-cos^2(x)) = 2 + cos(x)
2 - 2cos^2(x) = 2 + cos(x)
2cos^2(x) + cos(x) = 0
Factor:
cos(x)[2cos(x) + 1] = 0
cos(x) = 0 or 2cos(x)+1 = 0
x = 0 or cos(x) = -1/2
x = 0 or x = (2/3)pi or x = (4/3)pi
=====================================
Cheers,
Stan H.